# StableMaterials: Enhancing Diversity in Material Generation via Semi-Supervised Learning

### v3

@@ Abstract.
We detail the architecture and training process of StableMaterials, the integration of semi-supervised training within existing LDM frameworks. Comparative evaluations with state-of-the-art methods show the effectiveness of StableMaterials, highlighting its potential applications in computer graphics and beyond. StableMaterials will be made publicly available.
material appearance, generative models
journal: TOGccs: Computing methodologies Appearance and texture representations
Figure 1. We present StableMaterials, a diffusion–based model for materials generation through text or image prompting. Our approach enables high-resolution, tileable material maps, inferring both diffuse (Basecolor) and specular (Roughness, Metallic) properties, as well as the material mesostructure (Height, Normal).
[…]
Teaser Figure
1. Introduction
@@
Authoring of materials has been a long-standing challenge in computer graphics, requiring very specialized skills and a high level of expertise. To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [Deschaintre et al., 2018, 2019; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a], or generation from a set of conditions [Guehl et al., 2020; Guo et al., 2020; Hu et al., 2022; Zhou et al., 2022; He et al., 2023; Vecchio et al., 2024a, b]. While these approaches have reduced technical barriers to material creation, their effectiveness depends on the quality and diversity of training data, which can limit their use in real-world applications.
@@
Despite recent efforts to create large-scale materials datasets, such as Deschaintre et al. [2018], OpenSVBRDF [Ma et al., 2023], and MatSynth [Vecchio and Deschaintre, 2024], these datasets are limited in diversity [Zhou et al., 2023], not capturing the vast range observed in large-scale image datasets such as LAION [Schuhmann et al., 2022]. These limitations can constrain the capabilities of learning-based approaches, potentially creating gaps in their generative capabilities and affecting realism and diversity.
@@
Fine-tuning has become a common practice to reduce these gaps in training data by leveraging existing knowledge from large-scale pretrained models. Techniques like Low-Rank Adaptation (LoRA) [Hu et al., 2021] effectively fine-tune models while preventing catastrophic forgetting. Methods such as Diff-Instruct [Luo et al., 2024], on the other hand, employ distillation strategies to transfer knowledge from pretrained models. However, while fine-tuning or distillation within the same domain are straightforward, they pose significant challenges across different domains (e.g., image to material).
@@
To overcome these limitations, we introduce StableMaterials, an approach that takes advantage of semi-supervised adversarial training to: (1) include unannotated (non-PBR) samples in training, and (2) distill knowledge from a large-scale pretrained SDXL [Podell et al., 2023] model. In particular, we use a pretrained SDXL to generate unannotated texture samples from text prompts. However, since StableMaterials is trained to produce SVBRDF maps, we cannot perform direct supervision using the generated textures. To include these textures in the training of our methods, we learn a common latent representation between textures and materials; then, we complement the traditional supervised loss, with an unsupervised adversarial loss, forcing the model to also generate realistic maps for unannotated samples and close the gap between the two data distributions. In addition, drawing inspiration from the SDXL [Podell et al., 2023] architecture, we use a diffusion-based refinement model to enhance the visual quality of the samples and achieve high-resolution generation. We initially generate materials at the model base resolution of 512x512, and subsequently apply our refiner using SDEdit [Meng et al., 2021] and patched diffusion. This approach allows to achieve high resolution while constraining the patched generation, ensuring consistency and memory efficiency. Subsequently, we distill a latent consistency model [Song et al., 2023] that allows fast generation by reducing the number of inference steps to four steps per stage. However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [Vecchio et al., 2024a] to achieve tileability. To solve this issue, we propose a novel features rolling technique, which shifts the tensor rolling from the diffusion step to the U-Net architecture by directly shifting the feature maps within each convolutional and attention layer.
@@
  • StableMaterials, a new diffusion–based model for PBR material generation, leveraging a semi-supervised learning approach to incorporate unannotated data during training.
  • A novel distillation technique to bridge the gap with large-scale models by including unannotated data in the training.
  • A novel “features rolling” approach to tileability, minimizing the visual artifacts produced by fewer diffusion steps.
  • State-of-the-art performance in PBR materials generation.
2. Related Work
Figure 2. Architecture of StableMaterials. The base model generates a low resolution materials of size 512x512. This generation is then upscaled and refined using SDEdit [Meng et al., 2021] by the refiner model using a patched approach to limit memory requirements.
[…]
Method Figure
@@ 2. Related Work
Materials synthesis is an open challenge in computer graphics [Guarnera et al., 2016] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [Deschaintre et al., 2018; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Deschaintre et al., 2019; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a]. Building on the success of generative models, several approaches to materials generation have emerged, including Guehl et al. [2020] which combines procedural structure generation with data-driven color synthesis; MaterialGAN [Guo et al., 2020], a generative network which produces realistic SVBRDF parameter maps using the latent features learned from a StyleGAN2 [Karras et al., 2020a]; Hu et al. [2022] which generates new materials transferring the micro and meso-structure of a texture to a set of input material maps; and TileGen [Zhou et al., 2022], a generative model capable of producing tileable materials, conditioned by an input pattern but limited to class-specific training. Recent appraoches have focused on leveraging the generative capabilities of diffusion models for materials generation. In particular, Vecchio et al. [2024a] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image. MatFuse [Vecchio et al., 2024b] extends generation control with multimodal conditioning, and enables editing ofexisting materials via ‘volumetric’ inpainting. MaterialPalette [Lopes et al., 2024] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [Hu et al., 2021] for each picture. Substance 3D Sampler [Adobe, 2024] recently introduced a pipeline to generate materials by first synthesizing a texture via text conditioning. However, these methods often lack diversity, struggling with complex material representation, or depend on image generation models, requiring additional steps to estimate the material parameters. StableMaterials overcomes these limitations by including a wider variety of unannotated material samples via a semi-supervised training and improves on inference time by distilling a latent consistency model.
@@
Image generation is a long-standing challenge in computer vision, primarily due to the high dimensionality of images and complex data distributions. Generative Adversarial Networks (GAN) [Goodfellow et al., 2014] enabled the creation of high-quality images [Karras et al., 2017; Brock et al., 2018; Karras et al., 2020b], yet they suffer from unstable training [Arjovsky et al., 2017; Gulrajani et al., 2017; Mescheder, 2018], leading to mode collapse behaviour.
Diffusion Models (DMs) [Sohl-Dickstein et al., 2015; Ho et al., 2020], particularly the more efficient Latent Diffusion (LDM) architecture [Rombach et al., 2022], have emerged as an alternative to GANs, achieving state-of-the-art results in image generation tasks [Dhariwal and Nichol, 2021], besides showing a more stable training. Following the success of LDMs, research has focused on improving the generation quality [Podell et al., 2023] and reducing the number of inference steps to speed up the generation process [Song et al., 2023; Luo et al., 2023; Sauer et al., 2023b, 2024] through model distillation. Furthermore, due to the proliferation of large-scale pretrained models, several approaches were proposed to reuse their knowledge [Hu et al., 2021; Luo et al., 2024; Ruiz et al., 2023]. However, these approaches focus on fine-tuning within the same image domain, restricting their applicability to non-image domains.
3. Method
StableMaterials builds on MatFuse [Vecchio et al., 2024b], which adapts the LDM paradigm [Rombach et al., 2022] to synthesize high-quality, pixel-level reflectance properties for arbitrary materials (Fig. 2). We replace MatFuse’s multi-encoder VAE architecture (used to learn a disentangled latent representation of material maps) with a more resource-efficient single-encoder model, fine-tuned to preserve the same latent properties. In addition, we introduce a semi-supervised training strategy that distills knowledge from a large-scale SDXL model to increase generation diversity. Our method leverages latent consistency distillation [Song et al., 2023] and a novel feature rolling technique for fast, tileable generation. A dedicated refiner model enables high-resolution outputs while preserving global consistency.
3.1. Material Representation
StableMaterials generates SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [Cook and Torrance, 1982; Karis, 2013], using a GGX [Walter et al., 2007] distribution function, as well as the mesostructure of the material. Specifically, the generated maps include base color, normal, height, roughness, and metalness, where roughness specifies the specular lobe width and metalness indicates conductor regions.
3.2. Material Generation
The generative model consists of a compression VAE [Kingma and Welling, 2013] $\mathcal{E}$, encoding the material maps into a latent space, and a diffusion model [Rombach et al., 2022] $\epsilon_{\theta}$, modeling the distribution of these latent features.
@@ 3.2. Material Generation
We first train a multi-encoder VAE with encoders $\mathcal{E}={\mathcal{E}_{1},\dots,\mathcal{E}_{N}}$ and a decoder $\mathcal{D}$. Each encoder $\mathcal{E}_{i}$ encodes the $i^{th}$ map $\textbf{M}_{i}$ into a latent vector $z_{i}$, and the concatenated tensor $z=concat(z_{1},\dots,z_{N})$ is decoded to reconstruct the maps $\hat{M}=\mathcal{D}(z)$. Following [Vecchio et al., 2024b], training uses pixel-space $L_{2}$ loss, perceptual LPIPS [Zhang et al., 2021] loss, a patch-based adversarial loss [Isola et al., 2017; Dosovitskiy and Brox, 2016], and a rendering loss [Deschaintre et al., 2018], with a Kullback-Leibler penalty [Kingma and Welling, 2013; Rezende et al., 2014] to regularize the latent space.
Afterward, we fine-tune a single-encoder model. We freeze the original decoder and train only the encoder $\mathcal{E}^{\prime}$ to compress concatenated material maps into the same latent tensor $z$. This preserves the disentangled latent representation and maintains compression efficiency with fewer parameters, as shown in Sec. 4.4.1.
@@
We train a diffusion model on the compressed latent representation $z$ of the material. This model, based on Rombach et al. [2022], uses a time-conditional U-Net [Ronneberger et al., 2015] to denoise the latent vectors $z$. During training, we generate noised latent vectors using a deterministic forward diffusion process $q\left(z_{t}|z_{t-1}\right)$, as defined in Ho et al. [2020], transforming them into an isotropic Gaussian distribution. The diffusion network $\epsilon_{\theta}$ learns the backward diffusion $q(z_{t-1}|z_{t})$ to denoise and reconstruct the original latent vector. The model training is described in Sec. 3.3
@@
StableMaterials allows flexible control over the generated material via text or image prompts describing high-level appearance. We encode the text or image condition using a pretrained CLIP model [Radford et al., 2021], which outputs a single feature vector. To make training robust, we alternate between image and text prompts. Specifically, we use (i) an ambient-lit rendering of the material as an image condition and (ii) a text caption for text prompts; when no caption is available, we generate short descriptive tags as an alternative. In each training batch, one modality is randomly dropped—text with 75% probability, image with 25%—to balance the two conditions.
3.3. Semi-Supervised Adversarial Distillation
Figure 3. Semi-Supervised training. Both the SDXL model and StableMaterials are prompted to generate the same material. The supervised $\mathcal{L}_{2}$ loss, between the estimated noise and the added noise, is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
[…]
Method Figure
To bridge the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model. However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties) during training. Our method combines a supervised loss on annotated materials with an unsupervised adversarial loss that aligns the distribution of SDXL-generated textures with that of real materials in a shared latent space.
The supervised term ensures correct reconstruction of ground-truth material properties, while the adversarial term guides the model to learn from unannotated textures by treating them as partially labeled data. To accomplish this, we introduce a latent discriminator ($LD$) that distinguishes between real material features and generated ones, effectively forcing the generator to produce material-like features even from unannotated textures. Our training combines:
  1. (1)
    Supervised Loss ($\mathcal{L}{\text{sup}}$): Ensures reconstruction of material properties, maintaining physical plausibility.
  2. (2)
    Adversarial Loss ($\mathcal{L}{\text{adv}}$): Guides the generator to map both materials and textures to a shared feature distribution, enabling diversity while maintaining realism.
Our training strategy avoids mode collapse being primarely supervised, with the adversarial loss working as a distillation strategy.
Supervised Loss.
@@ 3.3. Semi-Supervised Adversarial Distillation
(1) $\mathcal{L}{\text{sup}}=\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon_{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{mat}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}+\alpha~{}\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{tex}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}.$
@@
(2) $\mathcal{L}{\text{adv}}=-\mathbb{E}{\mathbf{z}\sim p(\mathbf{z})}\Bigl{[}LD\bigl{(}z_{t-1}\bigr{)}\Bigr{]},$
@@
Latent Discriminator.
We follow Sauer et al. [2021, 2023a] and train $LD$ with a hinge loss [Lim and Ye, 2017], comparing real latent embeddings $z_{t-1,\text{real}}$ (encoded from the VAE) against fake latent embeddings $z_{t-1,\text{fake}}$ (denoised by the generator). The discriminator operates in a time-conditional fashion: it receives the same timestep $t$ and the CLIP embedding as the generator. Architecturally, $LD$ mirrors the U-Net encoder of the diffusion model and is initialized with the same weights, leveraging its understanding of the latent space to effectively guide the generator.
Unlike previous works that use adversarial distillation for fast generation [Sauer et al., 2023b, 2024], our approach bridges the domain gap between materials and textures. By training primarily with supervised terms while using adversarial guidance, the generator learns plausible material features even from unannotated textures. The adversarial component ensures SDXL-generated textures map to realistic material latents, effectively countering any shading artifacts and enriching the diversity of generated materials.
3.4. Fast High-Resolution Generation

### v4

@@ Abstract
We detail the architecture and training process of StableMaterials, the integration of semi-supervised training within existing LDM frameworks. Comparative evaluations with state-of-the-art methods show the effectiveness of StableMaterials, highlighting its potential applications in computer graphics and beyond. StableMaterials is publicly available at https://gvecchio.com/stablematerials.
1 Introduction
@@
Authoring of materials has been a long-standing challenge in computer graphics, requiring very specialized skills and a high level of expertise. To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [6, 7, 31, 32, 39, 16, 65, 59, 3, 11, 60], or generation from a set of conditions [14, 17, 23, 66, 18, 60, 61]. While these approaches have reduced technical barriers to material creation, their effectiveness depends on the quality and diversity of training data, which can limit their use in real-world applications.
@@
Despite recent efforts to create large-scale materials datasets, such as Deschaintre et al. [6], OpenSVBRDF [38], and MatSynth [58], these datasets are limited in diversity [67], not capturing the vast range observed in large-scale image datasets such as LAION [55]. These limitations can constrain the capabilities of learning-based approaches, potentially creating gaps in their generative capabilities and affecting realism and diversity.
@@
Fine-tuning has become a common practice to reduce these gaps in training data by leveraging existing knowledge from large-scale pretrained models. Techniques like Low-Rank Adaptation (LoRA) [22] effectively fine-tune models while preventing catastrophic forgetting. Methods such as Diff-Instruct [37], on the other hand, employ distillation strategies to transfer knowledge from pretrained models. However, while fine-tuning or distillation within the same domain are straightforward, they pose significant challenges across different domains (e.g., image to material).
@@
To overcome these limitations, we introduce StableMaterials, an approach that takes advantage of semi-supervised adversarial training to: (1) include unannotated (non-PBR) samples in training, and (2) distill knowledge from a large-scale pretrained SDXL [44] model. In particular, we use a pretrained SDXL to generate unannotated texture samples from text prompts. However, since StableMaterials is trained to produce SVBRDF maps, we cannot perform direct supervision using the generated textures. To include these textures in the training of our methods, we learn a common latent representation between textures and materials; then, we complement the traditional supervised loss, with an unsupervised adversarial loss, forcing the model to also generate realistic maps for unannotated samples and close the gap between the two data distributions. In addition, drawing inspiration from the SDXL [44] architecture, we use a diffusion-based refinement model to enhance the visual quality of the samples and achieve high-resolution generation. We initially generate materials at the model base resolution of 512x512, and subsequently apply our refiner using SDEdit [40] and patched diffusion. This approach allows to achieve high resolution while constraining the patched generation, ensuring consistency and memory efficiency. Subsequently, we distill a latent consistency model [57] that allows fast generation by reducing the number of inference steps to four steps per stage. However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [60] to achieve tileability. To solve this issue, we propose a novel features rolling technique, which shifts the tensor rolling from the diffusion step to the U-Net architecture by directly shifting the feature maps within each convolutional and attention layer.
@@
  • StableMaterials, a new diffusion–based model for PBR material generation, leveraging semi-supervised learning to incorporate unannotated data during training.
  • A novel adversarial distillation technique to bridge the gap with large-scale models.
  • A novel “features rolling” approach to tileability, minimizing the artifacts produced by fewer diffusion steps.
  • State-of-the-art performance in materials generation.
2 Related Work
Figure 1: Architecture of StableMaterials. The base model generates a low resolution materials of size 512x512. This generation is then upscaled and refined using SDEdit [40] by the refiner model using a patched approach to limit memory requirements.
@@ 2 Related Work
Materials synthesis is an open challenge in computer graphics [13] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [6, 31, 32, 39, 16, 65, 59, 7, 3, 11, 60]. Building on the success of generative models, several approaches to materials generation have emerged, including Guehl et al. [14] which combines procedural structure generation with data-driven color synthesis; MaterialGAN [17], a generative network which produces realistic SVBRDF parameter maps using the latent features learned from a StyleGAN2 [28]; Hu et al. [23] which generates new materials transferring the micro and meso-structure of a texture to a set of input material maps; and TileGen [66], a generative model capable of producing tileable materials, conditioned by an input pattern but limited to class-specific training. Recent appraoches have focused on leveraging the generative capabilities of diffusion models for materials generation. In particular, Vecchio et al. [60] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image. MatFuse [61] extends generation control with multimodal conditioning, and enables editing of existing materials via ‘volumetric’ inpainting. MaterialPalette [34] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [22] for each picture. Substance 3D Sampler [1] recently introduced a pipeline to generate materials by first synthesizing a texture via text conditioning. However, these methods often lack diversity, struggling with complex material representation, or depend on image generation models, requiring additional steps to estimate the material parameters. StableMaterials overcomes these limitations by including a wider variety of unannotated material samples via a semi-supervised training and improves on inference time by distilling a latent consistency model.
@@
Image generation is a long-standing challenge in computer vision, due to the high dimensionality of images and complex data distributions. Generative Adversarial Networks (GAN) [12] enabled the creation of high-quality images [26, 4, 27], yet they suffer from unstable training [2, 15, 42] and mode collapse behaviour.
Diffusion Models (DMs) [56, 21], particularly the more efficient Latent Diffusion (LDM) architecture [47], have emerged as an alternative to GANs, achieving state-of-the-art results in image generation tasks [8], besides showing a more stable training. Following the success of LDMs, research has focused on improving the generation quality [44] and reducing the number of inference steps to speed up the generation process [57, 36, 53, 54] through model distillation. Furthermore, due to the proliferation of large-scale pretrained models, several approaches were proposed to reuse their knowledge [22, 37, 49]. However, these approaches focus on fine-tuning within the same image domain, restricting their applicability to non-image domains.
3 Method
StableMaterials builds on MatFuse [61], which adapts the LDM paradigm [47] to synthesize high-quality, pixel-level reflectance properties for arbitrary materials (Fig. 1). We replace MatFuse’s multi-encoder VAE architecture (used to learn a disentangled latent representation of material maps) with a more resource-efficient single-encoder model. We introduce a semi-supervised training strategy that distills knowledge from a large-scale SDXL model to increase generation diversity, and leverage latent consistency distillation [57] and a novel feature rolling technique for fast, tileable generation. A dedicated refiner model enables high-resolution outputs while preserving global consistency.
3.1 Material Representation
StableMaterials generates SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [5, 25], using a GGX [62] distribution function, as well as the mesostructure of the material. Specifically, the generated maps include base color, normal, height, roughness, and metalness, where roughness specifies the specular lobe width and metalness indicates conductor regions.
3.2 Material Generation
The generative model consists of a compression VAE [30] $\mathcal{E}$, encoding the material maps into a latent space, and a diffusion model [47] $\epsilon_{\theta}$, modeling the distribution of these latent features.
@@ 3.2 Material Generation
We first train a multi-encoder VAE with encoders $\mathcal{E}={\mathcal{E}_{1},\dots,\mathcal{E}_{N}}$ and a decoder $\mathcal{D}$. Each encoder $\mathcal{E}_{i}$ encodes the $i^{th}$ map $\textbf{M}_{i}$ into a latent vector $z_{i}$, and the concatenated tensor $z=concat(z_{1},\dots,z_{N})$ is decoded to reconstruct the maps $\hat{M}=\mathcal{D}(z)$. Following [61], training uses pixel-space $L_{2}$ loss, perceptual LPIPS [64] loss, a patch-based adversarial loss [24, 9], and a rendering loss [6], with a Kullback-Leibler penalty [30, 46] to regularize the latent space.
Afterward, we fine-tune a single-encoder model. We freeze the original decoder and train only the encoder $\mathcal{E}^{\prime}$ to compress concatenated material maps into the same latent tensor $z$. This preserves the disentangled latent representation and maintains compression efficiency with fewer parameters, as shown in Sec. 4.4.
@@
We train a diffusion model on the compressed latent representation $z$ of the material. This model, based on Rombach et al. [47], uses a time-conditional U-Net [48] to denoise the latent vectors $z$. During training, we generate noised latent vectors using a deterministic forward diffusion process $q\left(z_{t}|z_{t-1}\right)$, as defined in Ho et al. [21], transforming them into an isotropic Gaussian distribution. The diffusion network $\epsilon_{\theta}$ learns the backward diffusion $q(z_{t-1}|z_{t})$ to denoise and reconstruct the original latent vector. The model training is described in Sec. 3.3
Figure 2: Semi-Supervised training. Both the SDXL model and StableMaterials are prompted to generate the same material. The supervised $\mathcal{L}_{2}$ loss, between the estimated noise and the added noise, is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
@@
StableMaterials allows flexible generation control via text or image prompts describing high-level appearance. We encode the text or image condition using a pretrained CLIP model [45], which outputs a single feature vector. To make training robust, we alternate between image and text prompts. Specifically, we use (i) an ambient-lit rendering of the material as an image condition and (ii) a text caption for text prompts; when no caption is available, we generate short descriptive tags as an alternative. In each training batch, one modality is randomly dropped—text with 75% probability, image with 25%—to balance the two conditions.
3.3 Semi-Supervised Adversarial Distillation
To bridge the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model. However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties) during training. Our method combines a supervised loss on annotated materials with an unsupervised adversarial loss that aligns the latent distribution of SDXL-generated textures with that of real materials.
The supervised Loss ($\mathcal{L}{\text{sup}}$) ensures correct reconstruction of ground-truth material properties, maintaining physical plausibility. The adversarial Loss ($\mathcal{L}{\text{adv}}$) guides the generator to map both materials and textures to a shared feature distribution, enabling diversity while maintaining realism. We introduce a latent discriminator ($LD$) that distinguishes between real material features and generated ones, effectively forcing the generator to produce material-like features even from unannotated textures. This strategy avoids mode collapse being primarely supervised, with the adversarial loss working as a distillation strategy.
Supervised Loss
@@ 3.3 Semi-Supervised Adversarial Distillation
$\mathcal{L}_{\text{sup}}$ $=\mathbb{E}_{t,z_{0},\epsilon}\left[\lVert\epsilon_{t}-\epsilon_{\theta}(z_{t,\text{mat}},t)\rVert^{2}\right]$ (1)
$\quad+\alpha\,\mathbb{E}_{t,z_{0},\epsilon}\left[\lVert\epsilon_{t}-\epsilon_{\theta}(z_{t,\text{tex}},t)\rVert^{2}\right].$ (2)
@@
$\mathcal{L}{\text{adv}}=-\mathbb{E}{\mathbf{z}\sim p(\mathbf{z})}\Bigl[LD\bigl(z_{t-1}\bigr)\Bigr],$ (3)
@@
Latent Discriminator
We follow Sauer et al. [51, 52] and train $LD$ with a hinge loss [33], comparing real latent embeddings $z_{t-1,\text{real}}$ (encoded from the VAE) against fake latent embeddings $z_{t-1,\text{fake}}$ (denoised by the generator). The discriminator operates in a time-conditional fashion: it receives the same timestep $t$ and the CLIP embedding as the generator. Architecturally, $LD$ mirrors the U-Net encoder of the diffusion model and is initialized with the same weights, leveraging its understanding of the latent space to effectively guide the generator. Only ground-truth material latents are treated as real samples by the discriminator.
Unlike previous works that use adversarial distillation for fast generation [53, 54], our approach bridges the domain gap between materials and textures. By training primarily with supervised terms while using adversarial guidance, the generator learns plausible material features even from unannotated textures. The adversarial component ensures SDXL-generated textures map to realistic material latents, effectively countering any shading artifacts and enriching the diversity of generated materials.
3.4 Fast High-Resolution Generation